OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 zero 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

@jim_thompson5910 can you help me plz

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

u know substitution method ??

OpenStudy (anonymous):

Yea

OpenStudy (gorv):

from first can u find value of y in term of x???

OpenStudy (gorv):

u there ??

OpenStudy (gorv):

add y on both side

OpenStudy (gorv):

-2y+2y=0

OpenStudy (gorv):

2x-y=6 move y on right hand side so its sign will change 2x=6+y 2x-6=y

OpenStudy (gorv):

now replace y with this value in second equation